ome Assume that you have a sample of n-5, with the sample mean X, -43, and a sample standard deviation of S, 4, and you have need answer to the second and third part
Answer #1

When the population variances are assumed to be equal, then the formula is:

t = \frac{\bar X_1 - \bar X_2}{\sqrt{ \frac{(n_1-1)s_1^2 + (n_2-1)s_2^2}{n_1+n_2-2}(\frac{1}{n_1}+\frac{1}{n_2}) } }

\rightarrow t = \frac{43 - 37}{\sqrt{ \frac{(5-1)*4^2 + (8-1)*5^2}{5+8-2}(\frac{1}{5}+\frac{1}{8}) } }

\rightarrow t = {\color{DarkRed} 2.26}

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!